Circular No. 7376

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 E1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports that several people browsing the SOHO Web site (including M. Meyer, M. Boschat, T. Harincar, and M. Oates) noted another Kreutz sungrazing comet in SOHO LASCO data. Observed in both the C2 and C3 telescope data, this object showed a tail of length ∼ 10′. Astrometry and parabolic orbital elements appear on MPEC 2000-E27.

2000 UT  α_{2000}  δ_{2000}
Mar. 4.404  23^{h}27.^{m}4  − 7^{o}27′

SUNGRAZING COMETS Biesecker also reports observations (measures by D. Hammer and him-

self, reductions by B. G. Marsden) of some older Kreutz sungrazing comets seen in SOHO C3 data; full astrometry and parabolic orbital elements appear on the MPECs indicated below. Comet C/1999 O2 developed a short tail, C/1999 O3 had a short tail evident, C/1999 Q3 showed a tail, and C/1999 R3 showed evidence for a tail; the other four comets showed no evident tail. Comets C/1999 O2 and C/1999 P4 were discovered by D. Lewis, C/1999 P5 by A. Vourlidas, C/1999 Q2 and C/1999 R3 by K. Schenk, and C/1999 Q3 by Biesecker, while comets C/1999 O3 and C/1999 Q1 were first noted by T. Lovejoy via the SOHO Web page.

1999 UT  α_{2000}  δ_{2000}  MPEC  Comet
July 21.888  7^{h}49.^{m}5 +18^{o}13′ 2000-E28  C/1999 O2
31.446  8 22.4 +16 25  2000-E29  C/1999 O3
Aug. 9.429  8 58.2 +14 39  2000-E30  C/1999 P4
12.138  9 06.3 +13 30  2000-E31  C/1999 P5
23.821  9 51.1 +10 39  2000-E32  C/1999 Q1
24.471  9 54.2 +10 11  2000-E33  C/1999 Q2
30.738 10 10.4 + 7 49  2000-E34  C/1999 Q3
Sept. 4.863 10 25.5 + 5 40  2000-E35  C/1999 R3
1999 LD_{31}

R. P. Binzel, Massachusetts Institute of Technology, reports that 0.5– 1.0-μm spectra of 1999 LD_{31}, obtained on Mar. 1.3 UT using the Kitt Peak 4-m reflector, display a constant reflectance spectral slope of 0.12 ± 0.02 per 100 nm (consistent with that of a typical D-type minor planet).

2000 March 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green